OpenStudy (anonymous):

whats the derivative of log base 7 to x?

OpenStudy (anonymous):

log of 'a' to any number b= loga/log b where both log a and log b are to the base 'e'. cnvert your sum into this form and then find the integral.

OpenStudy (anonymous):

\[ \log_7{x} = \frac{\ln(x)}{\ln(7)} \]